Suppose the life of a certain brand of vacuum cleaner is normally distributed with a mean life of 12 years and a standard deviation of 4 years. A vacuum cleaner is randomly chosen. For how many years should the manufacturer warranty its vacuum cleaners if it is willing to replace only 1% of the vacuum cleaners it sells.
Answer by stanbon(75887)

Suppose the life of a certain brand of vacuum cleaner is normally distributed with a mean life of 12 years and a standard deviation of 4 years. A vacuum cleaner is randomly chosen. For how many years should the manufacturer warranty its vacuum cleaners if it is willing to replace only 1% of the vacuum cleaners it sells.
----
Find the z-value with a right-tail of 1%.
invNorm(0.99) = 2.3263
---
Solve for "x":
x = zs+u
x = 2.3263*4+12 = 21.31 years (time to warranty the cleaners)
====================================
Cheers,
Stan H.